# Cohort Discovery: A Survey on LLM-Assisted Clinical Trial Recruitment

### v1

@@ Abstract
Recent advances in LLMs have greatly improved general-domain NLP tasks. Yet, their adoption in critical domains, such as clinical trial recruitment, remains limited. As trials are designed in natural language and patient data is represented as both structured and unstructured text, the task of matching trials and patients benefits from knowledge aggregation and reasoning abilities of LLMs. Classical approaches are trial-specific and LLMs with their ability to consolidate distributed knowledge hold the potential to build a more general solution. Yet recent applications of LLM-assisted methods rely on proprietary models and weak evaluation benchmarks. In this survey, we are the first to analyze the task of trial-patient matching and contextualize emerging LLM-based approaches in clinical trial recruitment. We critically examine existing benchmarks, approaches and evaluation frameworks, the challenges to adopting LLM technologies in clinical research and exciting future directions.
Cohort Discovery: A Survey on LLM-Assisted Clinical Trial Recruitment
@@ 1 Introduction
Clinical trials evaluate the effects of an intervention on human health. Selecting the precise and required size of patient population is crucial for trial completion. According to various estimates, more than 50% of aborted clinical trials fail due to low accrual rates, and 80% of all clinical trials do not manage to recruit the required patient cohorts within the allotted time Clinical Trials Arena (2012); Williams et al. (2015); Pharmaceutical Technology (2019). While the trend has steadily declined over the past decade, with the intensive use of technology-aided solutions, efficient patient recruitment remains the most crucial bottleneck in clinical trial research Clinical Trials Arena (2022), thereby hampering the rate at which potentially life-saving therapies or new vaccines can be introduced.
Figure 1: Steps in a typical patient recruitment process vs. our proposed flow with expert interaction.
Clinical trials have a dual nature, consisting of both trial-agnostic and trial-specific requirements, making it challenging to design generalized approaches. This has traditionally resulted in highly specialized matching processes that work for a particular trial or patient population. The resulting patient recruitment process is linear, as illustrated (in orange) in Figure 1, leaving little room for expert intervention and feedback. It starts with extracting information from a trial description and available patient records, then matching explicit trial eligibility criteria and patient data to determine patient eligibility. Finally, experts review the selected patients before contacting them.
An expert reviews several hundred patients Ni et al. (2015) and can end up spending hours on one patient, incurring significant costs Penberthy et al. (2012). In their study on five trials, Penberthy et al. (2010) showed that automated matching, via table queries and string-matching algorithms, saved between 165 hours to 1,329 hours of reviewing time when compared to manual evaluation. As more patient data becomes digitally available, research into applying machine intelligence and developing explainable systems for clinical researchers to correctly interpret model predictions has gained traction Murdoch and Detsky (2013); Payrovnaziri et al. (2020); von Itzstein et al. (2021).
Generative LLMs can serve as knowledge aggregators and streamline clinical trial design and trial-patient matching (in purple in Figure 1). However, the clinical trial recruitment process has seen relatively low adoption of transformer language models He et al. (2020); Harrer (2023); Lu et al. (2024). Instruction fine-tuned LLMs for generating explainable trial-patient matches has received attention only since the past one year Jin et al. (2024); Nievas et al. (2024); Rybinski et al. (2024); Wornow et al. (2025). Given the sensitive nature of clinical trials and the rapidly evolving landscape of LLM technology, there is no prior work surveying this area. We fill this gap by critically examining the trial-patient matching approaches, the method used, the benchmarks available and the evaluation strategies. The existing systematic reviews of NLP methods and their performance Kim and Quintana (2022); Idnay et al. (2022); Kantor and Morzy (2024) are desiged to follow strict inclusion/exclusion criteria and search strategies Khan et al. (2003); Moher et al. (2009). This practice, essential in medical research for measuring the association between intervention and outcomes, misses the larger picture of contextualizing LLMs in clinical trial recruitment settings. In contrast, this work follows the perspective of surveys in computer science (see Appendix B) to provide a more comprehensive view of the role of LLMs in clinical trial recruitment.
@@
Cohort discovery has been a long-standing problem and LLM-assisted methods to improve trial-patient matching has only recently gained momentum. In this work, we formalize the problem of trial-patient matching, and differentiate between two major approaches (trial-centric and patient-centric), the associated tasks (classification vs. ranking) and their effect on the choice of benchmarks (longitudinal vs. short patient descriptions, single vs. multi-trial). We identify the important bottlenecks in data availability and task formalization. We introduce a taxonomy of errors in LLM-generations for trial-patient matching. Finally, we provide actionable steps for the community on four promising directions.
@@ 2 Background
Also known as patient recruitment/enrollment/(pre-)screening, it is the process of finding relevant patients (or cohort) for a clinical trial. The steps in patient recruitment, as shown in Figure 1, include data sourcing, information extraction, matching, and expert review.
Eligibility Criteria
They are explicit conditions for (inclusion) or against (exclusion) enrollment, usually expressed in natural language and designed by trial investigators to select cohorts. These conditions have varying difficulty levels, ranging from simple tabular operations (age or gender requirements) to semantic matching (medical condition or medications) and inference (reasoning with negation, patient history).
Patient Records
They refer to any form of data, in a combination of structured and unstructured formats, used to describe a patient. When digitized, it is often called an Electronic Health Record (EHR). This data includes the patient’s demographic information (age, gender, ethnicity), clinical notes, medications, billing history, and test results, that are collected over a period of years.
@@
Transformer language models have greatly improved information retrieval and extraction tasks. There exist specialized resources, such as, PubmedQA Jin et al. (2019), MIMIC-III & IV Johnson et al. (2016, 2020) and BioRED Luo et al. (2022), to train encoder language models Radford (2018); Devlin et al. (2019); Raffel et al. (2020) on medical domain (e.g., BioBERT Lee et al. (2020), PubmedBERT Gu et al. (2021), Criteria2Query2.0 Fang et al. (2022), MedCPT Jin et al. (2023)). Yet, Dorfner et al. (2024) show that fine-tuning generative LLMs on biomedical domain has limited relative performance gain. In fact, Alber et al. (2025) show that medical misinformation encountered during the fine-tuning affect downstream tasks. The Journal of American Medical Informatics Association (JAMIA), which published 41 articles on biomedical health and LLMs in a focus issue, observed a shift from domain-specific small models towards larger general-purpose LLMs Lu et al. (2024). Even among LLMs, the OpenAI Achiam et al. (2023) family of proprietary models (GPT3.5 and GPT4 being the most common) is used far more often than open-sourced LLaMA Touvron et al. (2023) or Mistral Jiang et al. (2023) models. Finally, Bedi et al. (2025) report that of more than 519 biomedical studies using LLMs, less than 5% used real patient data, tasks like summarization are understudied and fairness, bias, uncertainty and deployment considerations are rarely measured.
@@
Historically, patient recruitment has been a manual process with minimal automation via pre-defined database queries Thadani et al. (2009), and, more recently, automatic translations of free text into database queries Yuan et al. (2019); Fang et al. (2022); Park et al. (2024). A recent review reported low adoption rates of AI, particularly NLP, in clinical trials Kantor and Morzy (2024). Another review on understanding the role of NLP systems in patient recruitment found just 11 studies that evaluated NLP systems in clinical settings Idnay et al. (2022). Heterogeneous outcomes, diverse results and a dependence on small retrospective data drive the gap in NLP research and their adoption in real-world settings Idnay et al. (2024). A study by Idnay et al. (2023), investigating how clinical researchers screened patients, highlights the challenges of universal and domain-specific nature of the eligibility criteria and makes recommendations to build interactive, flexible and transparent recruitment strategies. Interestingly, when Corbaux et al. (2024) categorize trial matching tools for oncological trials as manual, semi-automatic and automatic, automatic methods still fall in the research and development phase, yet to be commercially available.
Review Methodology
We searched for relavant papers since 2019, using the keywords “clinical trial”, “cohort discovery”, “patient recruitment”, “trial recruitment”, “trial matching” in conjunction with “llm”, “language model”, “gpt”. We included method and dataset papers using LLMs and excluded papers that used LLMs in biomedical NLP (broader scope) or other health care applications (not clinical trials). Starting with the citations in the LLM papers we tracked the pre-LLM literature by recursively tracing their citations in newer publications. Since the topic “clinical trial recruitment” is vast, this approach efficiently captured the evolution of the predictive methods in this field.
3 Trial-Patient Matching
Trial-patient matching assigns a trial-patient pair one of the three labels of eligible, ineligible and irrelevant based on the patient record and eligibility criteria. Importantly, a trial-centric approach, taken by a trial investigator, matches patients from a hospital database to a specific trial. Alternatively, a patient-centric approach, taken by a patient or their healthcare provider, matches relevant trials from a trial registry to a patient.
This induces a priority, such that, a patient satisfying any exclusion criteria becomes ineligible, regardless of inclusion criteria matches. In Section 6, we elaborate on the lack of a formal task definition and its effects on trial-centric matching.
Task Benchmark Size Eval. Metrics Best Score Cohort
@@ 1.0 MRR
TREC CT Track 2022
@@ 0.72 MRR
TREC CT Track 2023
@@ 0.78 MRR
Best scores are aggregated from the Appendix of the proceedings of TREC 2021. ∗∗Borrowed from Rybinski et al. (2024) as TREC CT 2023 does not have a published track overview.
Table 1: Overview of public evaluation benchmarks for trial-patient matches covered in this survey.
4 Benchmarks and Corpora
We analyze five public benchmarks in context of the corpora and their evaluation metrics.
  • 2018 N2C2 Cohort Selection for Clinical Trials Stubbs et al. (2019). [^1] Currently unavailable as of 2024 Nov 6.
  • Trial ranking Koopman and Zuccon (2016).
@@ 4.1 Trial Data
ClinicalTrials.gov is one of the largest online database of clinical studies, including clinical trials and other observational studies. As of November 2024, it lists more than 390,000 clinical trials, thousands of which are active. These studies are submitted by investigators from over 200 countries. The euclinicaltrials.eu is another online registry for trials conducted in the European Union. It comprises more than 50,000 trials, with more than 7000 being active. Every trial comprises a study description, eligibility criteria and study plan among other details. The ClinicalTrials.gov database forms the primary source of trials in many academic studies.
As we will see in Section 5, earlier trial-specific approaches manually separated discrete data from free-text search. However, this separation is difficult to scale, and newer approaches focus on semantic matching. The N2C2 cohort selection benchmark is based on a single trial. Hence, systems specifically model the trial criteria. The TREC CT benchmarks source the ClinicalTrials.gov database for the task of trial ranking. However, additional annotation of the eligibility criteria denoting their type or complexity is unavailable.
@@ 4.2 Patient Data
Since patient data contains sensitive information, additional measures are required to remove personal information in order to release them as public repositories. The MIMIC database Johnson et al. (2016, 2020) is the largest public database of patient information. The latest version, MIMIC-IV, contains rich data on hospital admissions, medications, demographic data and notes of 299,712 de-identified patients. This dataset is not used in any of the five benchmarks. One reason might be a significant challenge in obtaining confirmed trial-patient match labels, which is easily available retrospectively from existing hospital records. The N2C2 cohort selection benchmark Stubbs et al. (2019), contains 288 de-identified longitudinal records of patients. The TREC CT benchmarks use synthetic patient profiles, comprising short textual patient descriptions and keywords defining cohorts.
4.3 Evaluation Metrics
Standard IR metrics (precision, recall, F1) and their derivatives (precision at rank (P@k), normalized discounted cumulative gains (NDCG@k), R-precision, mean reciprocal rank (MRR) and adaptive precision Moffat et al. (2015)) are the most popular evaluation metrics. Accuracy, specificity, and workload are also prevelant in the medical community. Notably, researchers in the medical domain have stressed the lack of definition and consistency of ground truth data to evaluate NLP tools in clinical trial matching Kim and Quintana (2022). The trial-patient ground-truth labels are obtained manually either from medical experts Koopman and Zuccon (2016); Soboroff (2021); Roberts et al. (2022) or retrospectively from past trials Stubbs et al. (2019). The TREC CT annotations were created from the top-k results pooled among all participating teams, resulting in tens of thousands of relevance judgments. Out of a total of 35,394 pooled trial-patient matches in the 2022 edition, 11% were judged as eligible, 9% as ineligible and 80% as not relevant with an average of 700 trials judged per patient Roberts et al. (2022). In the 2023 edition, the relevance judgments were more balanced between the three labels Rybinski et al. (2024). Measurement of human preference of LLM generations, bias in model predictions and the informativeness of LLM explanations ranges from limited to non-existent in the current works.
5 Trial-Patient Matching Methods
We broadly categorize systems into classical and LLM-based approaches and further group them by their direction of approach, trial/patient-centric. The LLM-based approaches focus on tasks, such as, criterion/trial-level predictions and relevance ranking, which in principle is direction-agnostic. Appendix A lists the works cited in this paper.
5.1 Classical Approaches
Trial-centric Matches
Majority of classical models use this approach build trial or cohort specific models. They follow a two-step method, starting with manually-designed rule-based discrete data filters, followed by sub-word matching Penberthy et al. (2010), vector matching Ni et al. (2015) or machine-learned classifiers Tun et al. (2023) for matching eligibility criteria or cohort keywords. Ni et al. (2015) showed a retrospective workload reduction to 50 patients down from 163 per trial. IBM Watson for clinical trial matching (IBM-CTM) is a proprietary commercial tool for trial-patient matching, which uses neural NLP techniques to match eligibility criteria to patients Helgeson et al. (2018). While one study on 3 trials measured a reduction in workload from 110 minutes to 86 minutes for screening 90 patients Beck et al. (2017), another study on 10 clinical trials and 102 patients evaluated the agreement in matches between IBM-CTM and expert annotation to be above 80% Alexander et al. (2020). Another line of work that focuses on defining a cohort by translating natural language criteria into formal query language for instantly querying patient databases Yuan et al. (2019).
Patient-centric Matches
Ni et al. (2015) create bag-of-words vectors to retrieve top trial matches for a patient and show a retrospective workload reduction (4 vs. 42 trials per patient) and 33% precision on a subset of 127 patients. The workload increases to 20 trials per patient and the precision drops to less than 4% when generalized to 215 patients. Zhang and Demner-Fushman (2017) train an SVM classifier on the top unigram and bigram features to predict if a trial is eligible for a cohort (e.g., HIV patients). Additional features derived from NER and lemmatization yielded small gains in performance.

### v2

# A Survey on LLM-Assisted Clinical Trial Recruitment

@@ Abstract
Clinical trials are designed in natural language and the task of matching them to patients, represented via both structured and unstructured textual data, benefits from knowledge aggregation and reasoning abilities of LLMs. LLMs with their ability to consolidate distributed knowledge hold the potential to build a more general solution than classical approaches that employ trial-specific heuristics. Yet, adoption of LLMs in critical domains, such as clinical research, comes with many challenges, such as, the size of the benchmarks, the dimensions of evaluation and data sensitivity. In this survey, we contextualize emerging LLM-based approaches in clinical trial recruitment. We examine the main components of the clinical trial recruitment process, discuss existing challenges in adopting LLM technologies in clinical research and exciting future directions.
A Survey on LLM-Assisted Clinical Trial Recruitment
@@ 1 Introduction
Clinical trials evaluate the effects of an intervention on human health. Selecting the precise and required size of patient population is crucial for trial completion. According to various estimates, more than 50% of aborted clinical trials fail due to low accrual rates, and 80% of all clinical trials do not manage to recruit the required patient cohorts within the allotted time Clinical Trials Arena (2012); Williams et al. (2015); Pharmaceutical Technology (2019). While the trend has steadily declined over the past decade, with the intensive use of technology-aided solutions, efficient patient recruitment remains the most crucial bottleneck in clinical trial research Clinical Trials Arena (2022). As electronic health records (EHRs) of patients become more accessible, clinical researchers adopt machine intelligence and develop explainable systems to correctly interpret model predictions Murdoch and Detsky (2013); Payrovnaziri et al. (2020); von Itzstein et al. (2021).
There are methods leveraging LLMs in clinical trial research for cohort retrieval and modelling Fang et al. (2022); Tian et al. (2023); Park et al. (2024); Liu et al. (2025a); Wang et al. (2025), trial design Reinisch et al. (2024); Curran et al. (2024); Bornet et al. (2025); Neehal et al. (2025), trial search White et al. (2023); Rybinski et al. (2020), trial matching Jin et al. (2024); Nievas et al. (2024); Wornow et al. (2025), predicting trial outcomes and duration Reinisch et al. (2024); Yue et al. (2024a, b); Liu et al. (2025c), assessing risk of bias Lai et al. (2024); Ji et al. (2025), and extracting clinical trial results Lee et al. (2024), even as the community is coming up with recommended practices for responsible use of AI in the entire drug development process Geraci et al. (2025).
Figure 1: Components in a patient recruitment process: conventional linear flow (in orange) vs. our proposed LLM-assisted interactive flow (in purple).
Figure 1 shows the components in clinical trial recruitment, namely, data sourcing, information extraction, matching, and evaluation. An expert reviews several hundred patients per trial and can end up spending hours on one patient, hence incurring significant costs Penberthy et al. (2012); Ni et al. (2015). Even simple automation using table queries and lexical searches saves between 165 hours to 1,329 hours of reviewing time when compared to manual evaluation Penberthy et al. (2010). In the past, the patient recruitment process has seen relatively low adoption of the pretrained language models He et al. (2020); Harrer (2023); Lu et al. (2024). Generative LLMs serve as knowledge aggregators, and through their reasoning and instruction-following capabilities, they have revived research in the task of trial and patient matching Jin et al. (2024); Nievas et al. (2024); Rybinski et al. (2024); Wornow et al. (2025).
Difference to Prior Work
Despite the rapidly evolving landscape of LLM technology, there is no prior work surveying this area. Gueguen et al. (2025) evaluate public trial matching tools and Layne et al. (2025) compare the efficacy of open and proprietary LLM-assisted trial-patient matching in oncology. Systematic reviews in this topic, bound by strict selection criteria and highly specific research question, do not to capture the broad perspectives Kim and Quintana (2022); Idnay et al. (2022); Kantor and Morzy (2024b); Chen et al. (2025) (Appendix A differentiates between systematic reviews and surveys). Although it provides an overview of LLM approaches used in clinical trial matching, Chen et al. (2025) only briefly discusses associated challenges and future work.
@@
We examine the main components in a trial recruitment process, presented in Figure 1. We formalize the problem of trial-patient matching. We analyze existing approaches via the tasks (classification vs. ranking), the directionality (trial-centric and patient-centric) and the choice of benchmarks (longitudinal vs. short patient descriptions, single vs. multi-trial). We present a taxonomy of errors for consistent evaluation of LLM-generated responses. We discuss the critical challenges associated with using LLMs in trial recruitment research. Finally, we present actionable steps towards an interactive patient recruitment (illustrated in purple in Figure 1).
@@ 2 Background
Also known as patient recruitment/enrollment/(pre-)screening, it is the process of matching patients (or a cohort) to a clinical trial via its eligibility criteria. Clinical trials have a dual nature, consisting of both universal and trial-specific requirements, making it challenging to design generalized approaches Idnay et al. (2023). This has traditionally resulted in linear trial-centric matching processes (illustrated in orange in Figure 1) with limited scope for interaction and feedback. Such systems generally have an initial data filter on the structured EHRs followed by keyword matches and concept identification Penberthy et al. (2010); Tun et al. (2023) or cohort-specific classifiers Zhang and Demner-Fushman (2017). Ni et al. (2015) represented both trials and patients as feature vectors, thus supporting retrievals for both trial and patient.
@@
The biomedical NLP landscape is shifting from specialized pretrained language models, such as, BioBERT Lee et al. (2020), BioLM Lewis et al. (2020), PubmedBERT Gu et al. (2021), BioGPT Luo et al. (2022), MedCPT Jin et al. (2023), among many others Wang et al. (2023), towards instruction-following and chat-enabled LLMs (commonly termed as generative AI), used as is Nori et al. (2023); Kung et al. (2023) or fine-tuned for domain alignment, such as, Med-PaLM Singhal et al. (2023), Med-Alpaca Han et al. (2023) and LLaVA-Med Li et al. (2023). We point our reader to Thirunavukarasu et al. (2023); Liu et al. (2025b) for further reading. The Journal of American Medical Informatics Association (JAMIA), which published 41 articles on biomedical health and LLMs in a focus issue, also observed this shift towards generative AI Lu et al. (2024). Among these, the OpenAI Achiam et al. (2023) family of proprietary models (GPT3.5 and GPT4 being the most common) is used far more often than open-sourced models Touvron et al. (2023); Bai et al. (2023); Jiang et al. (2023) models. Further, fine-tuning generative LLMs on the biomedical domain offers limited performance gain Dorfner et al. (2024). In fact, Alber et al. (2025) show that models are more prone to propagate medical misinformation encountered during fine-tuning despite performing well on benchmarks. Finally, Bedi et al. (2025) report that real patient data is used in less than 5% of the 519 biomedical studies using LLMs, and fairness, bias, uncertainty and deployment considerations are rarely assessed.
@@
Kantor and Morzy (2024b)’s study on AI adoption rates in eligibility criteria parsing, report very low adoption rates of generative AI, with BERT family of models being most popular and generative models coming into the picture only since 2024. A systematic review of the role of NLP systems in patient recruitment identified only 11 studies Idnay et al. (2022). Heterogeneous outcomes, diverse results, a dependence on small retrospective data and a lack of common standardized benchmarks drive the gap in NLP research and their adoption in real-world settings Idnay et al. (2024); Kantor and Morzy (2024b). A study by Idnay et al. (2023), investigating how clinical researchers screened patients, highlights the challenges of universal and domain-specific nature of the eligibility criteria and makes recommendations to build interactive, flexible and transparent recruitment strategies. Interestingly, when Corbaux et al. (2024) categorize tools for oncological trial matching, they indicate that the automatic methods still fall in the research and development phase, yet to be commercially available.
3 Methodology
We queried Google Scholar with the keywords “clinical trial”, “cohort discovery”, “patient recruitment”, “trial recruitment”, “trial matching” in conjunction with “llm”, “language model”, “gpt” and inspected the first ten result pages, between the years 2019 and 2025 (both inclusive). We also used Jin et al. (2024) and Wornow et al. (2025) as seeds and recursively traced their citations to efficiently captured the evolution of the predictive methods in this field. We organized papers that described LLM-based approaches and corpora for clinical trial recruitment into data sourcing (Section 2), information extraction and parsing (Section 5), trial-patient matching (Section 6) and evaluation (Section 7). Via our discussion on critical limitations in Section 8 and on promising directions towards an interactive patient recruitment in Section 9, we provide further contextualize LLMs in healthcare settings and other stages of the drug development process that directly effect on patient recruitment.
Task Benchmark Size Eval. Metrics Best Score #P LLM Usage (#P) Cohort
@@ 1.0 MRR
26 BERT-based keyword extraction (9) / query summarization (1),
Transformer-based rankers (10)
TREC CT 2022
@@ 0.72 MRR
12 Query reformulation using BERT / sequence-to-sequence models (3),
Transformer-based rankers (3)
TREC CT 2023
@@ 0.78 MRR
11 Query reformulation using LLMs (5),
Transformer-based rankers (6),
LLM prompt-based relevance prediction (4)
Best scores are aggregated from the Appendix of the proceedings of TREC 2021. No participants, since this was not a challenge. ∗∗Borrowed from Rybinski et al. (2024) as TREC CT 2023 does not have a published track overview.
Table 1: Overview of the public trial-patient matching benchmarks. #P is the number of participating teams. LLM usage (#P) tracks the number of participants using LLMs that we could verify from the proceedings.
4 Data Sourcing: Public Benchmarks
We analyze five trial-patient matching benchmarks.
  • 2018 N2C2 Cohort Selection Stubbs et al. (2019) for criterion-level eligibility prediction. [^1] Currently unavailable as of 2024 Nov 6.
  • Koopman and Zuccon (2016) for ranking trials.
@@ 4.1 Trial Data
The TREC CT benchmarks and Koopman and Zuccon (2016) source the ClinicalTrials.gov registry. Meanwhile, the N2C2 benchmark focuses on a single trial. ClinicalTrials.gov is one of the largest online databases of clinical studies submitted by investigators from over 200 countries. As of June 2025, it lists more than 400,000 clinical trials, thousands of which are active. The euclinicaltrials.eu is another online registry comprising over 50,000 trials from the European Union of which, 7000 are active. Every trial comprises a study description, eligibility criteria and study plan among other details.
@@ 4.2 Patient Data
The N2C2 benchmark contains 288 de-identified longitudinal records of patients. The trial ranking benchmarks use up to 75 synthetic patient profiles, comprising keywords defining cohorts Koopman and Zuccon (2016), short textual patient descriptions (TREC CT 2021, 2022) and questionnaire templates (TREC CT 2023). The MIMIC database Johnson et al. (2016, 2020) is the largest anonymized public database of structured and unstructured data of 299,712 patients. While this dataset is popular for training biomedical LLMs (as mentioned in Section 2), it is not used in any of the five benchmarks, possibly due to significant challenges in obtaining eligibility labels at this scale.
4.3 Annotated Labels
Two medical experts annotated 3,744 criterion-level labels in the N2C2 benchmark. The TREC CT annotations were created by pooling the top-k results from all participating teams. Medical experts manually annotated this pool of trial-patient matches Koopman and Zuccon (2016); Soboroff (2021); Roberts et al. (2022). Out of a total of 35,394 pooled trial-patient matches in the 2022 edition, 11% were judged as eligible, 9% as ineligible and 80% as not relevant with an average of 700 trials judged per patient Roberts et al. (2022). The relevance judgments were more balanced between the three labels Rybinski et al. (2024) in TREC CT 2023. All benchmarks depend on manual annotation from experts, which is time-consuming and challenging to scale Kim and Quintana (2022).
5 Information Extraction and Parsing
Extraction of medical entities evolved from a combination of rule-based heuristics and feature-based supervised sequence labelling models Kang et al. (2017); Yuan et al. (2019) via embedding-based neural models Khan et al. (2019); Tseo et al. (2020) to transformer-based pretrained models and LLMs Liu et al. (2021c); Zeng et al. (2020); Tian et al. (2021); Li et al. (2022); Murcia et al. (2024); Kantor and Morzy (2024a). Datta et al. (2024) use disease-specific prompting to extract structured information about entities and its attributes from criteria text. Gao et al. (2020); Zhang et al. (2020); Theodorou et al. (2023) use BERT-based embeddings to encode eligibility criteria and patient data. Patient data converted to search queries, via reformulation and expansion using LLMs, are particularly effective in trial retrieval Peikos et al. (2023); Rybinski et al. (2024); Peikos et al. (2024); Jin et al. (2024). Yuan et al. (2019); Tian et al. (2023); Park et al. (2024); Mugambi et al. (2024); Ziletti and D’Ambrosi (2025) use semantic parsing to translate eligibility criteria into logical forms ready for querying patient databases.
6 Trial-Patient Matching
6.1 Formalization
This induces a priority, such that, a patient satisfying any exclusion criteria becomes ineligible, regardless of inclusion criteria matches. Section 6.2 explores the task at varying levels of granularity, starting from criterion-level via trial-level predictions to trial ranking. In Section 8, we elaborate on the importance of formalization and its effects on trial-level aggregation.
Despite the task being direction-agnostic, there exist two directional approaches to tackle the matching problem, mainly due to data availability. First is the trial-centric approach, taken by a trial investigator, that matches longitudinal patient records to a specific trial. The 2018 N2C2 cohort selection is a trial-centric benchmark with criterion-level predictions. Alternatively, a patient-centric approach, taken by a patient or their healthcare provider, matches relevant trials from a trial registry to a short description of the patient. Koopman and Zuccon (2016) and the TREC CT 2021, 2022, and 2023 are patient-centric benchmarks for ranking trials.
Classical LLM-based ! Direction-specific approaches, applicable to a set of trials or a cohort * Direction-agnostic criteria- & trial-level prediction ! Trial-specific heuristics: filters and features * Generalizable across trials ! Evaluated on private patient data * Public benchmarks more common
Table 2: Differences between classical and LLM-based trial matching approaches.